Can someone help me with this word problem please..
Using the compound interest formula A + P(1 + i)^(n), where P is the original value of an investment, I is the interest rate per compounding period, n is the total number of compounding periods, and A is the value of the investment after n periods.
To save for college tuition, the parents of a preschooler invest $5000 in a bond fun that earns 6% annual interest compounded monthly. In approximately how many years will the investment be worth $15000?
Thank you!

Answer by stanbon(75887)

To save for college tuition, the parents of a preschooler invest $5000 in a bond fun that earns 6% annual interest compounded monthly. In approximately how many years will the investment be worth $15000?
-----------------
Correct Formula: A(t) = P(1+(r/n))^(nt)
---
15000 = 5000*(1+(0.06/12))^(12t)
3 = (1+ 0.005)^(12t)
Take the log of both sides to get:
12t*log(1.005) = log(3)
12t = [log(3)]/[log(1.005)]
12t = 220.2713
t = 18.36 years
========================
Cheers,
Stan H.